# From Code to Correctness: Closing the Last Mile of Code Generation with Hierarchical Debugging

### v2

@@ Abstract
While large language models have made significant strides in code generation, the pass rate of the generated code is bottlenecked on subtle errors, often requiring human intervention to pass tests, especially for complex problems. Existing LLM-based debugging systems treat generated programs as monolithic units, failing to address bugs at multiple levels of granularity, from low-level syntax errors to high-level algorithmic flaws. In this paper, we introduce Multi-Granularity Debugger (MGDebugger), a hierarchical code debugger by isolating, identifying, and resolving bugs at various levels of granularity. MGDebugger decomposes problematic code into a hierarchical tree structure of subfunctions, with each level representing a particular granularity of error. During debugging, it analyzes each subfunction and iteratively resolves bugs in a bottom-up manner. To effectively test each subfunction, we propose an LLM-simulated Python executor, which traces code execution and tracks important variable states to pinpoint errors accurately. Extensive experiments demonstrate that MGDebugger outperforms existing debugging systems, achieving an 18.9% improvement in accuracy over seed generations in HumanEval and a 97.6% repair success rate in HumanEvalFix. Furthermore, MGDebugger effectively fixes bugs across different categories and difficulty levels, demonstrating its robustness and effectiveness. [^1] Code and data available at https://github.com/YerbaPage/MGDebugger
1 Introduction
@@
Large language models (LLMs) such as GPT-4 (OpenAI, 2023), LLaMA (Touvron et al., 2023), and DeepSeek-Coder (Zhu et al., 2024) have made significant advances in AI-assisted coding tasks (Chen et al., 2021; Lu et al., 2021; Li et al., 2022). Trained on vast corpora of text and code, LLMs can understand and generate code snippets for various programming tasks, ranging from simple data structures to complex algorithmic problems (Li et al., 2022). These models have demonstrated proficiency in tasks such as code completion, bug detection, and even tackling competitive programming challenges.
@@
While the code generated by large models generally meets the requirements, it often contains critical errors that require human intervention to pass tests (Liu et al., 2023b; Dou et al., 2024). This has gradually led to a new development paradigm: large models generate the code, while humans fix it. Therefore, the “last mile”, as well as the most crucial step, of code generation is how to efficiently repair the code generated by large models.
@@
Numerous efforts have been made to debug LLM-generated code. The most popular way is to reuse the LLM generator to debug the generated code with the feedback from test case execution (Chen et al., 2023b; Zhong et al., 2024; Hu et al., 2024). While these methods increase the pass rates, they treat the erroneous program as a holistic set of statements (Chen et al., 2023b; Shinn et al., 2023; Zhong et al., 2024; Ding et al., 2024) regardless of the varying types and levels of failures. Failures of test cases arise from different levels of factors, from low-level syntactic errors to high-level algorithmic flaws. A holistic treatment overlooks the internal structure of the code and limits the effectiveness of the debugging systems, especially when dealing with complex programs that need debugging across different modules (Zeller, 2009; Tian et al., 2024).
@@
Extensive experiments with three models across three benchmarks demonstrate that MGDebugger significantly outperforms existing debugging methods, elevating accuracy from 75.6% to 94.5% on HumanEval (Chen et al., 2021) and achieving a remarkable 97.6% repair success rate on HumanEvalFix (Muennighoff et al., 2023). Ablation studies confirm the vital role of the hierarchical debugging strategy. We also evaluate MGDebugger’s effectiveness in handling diverse bug types and varying code lengths, highlighting its robustness and adaptability in real-world coding scenarios. Overall, these results underscore MGDebugger’s potential for enhancing the reliability of LLM-generated code.
2 Related Work
Code Generation with LLMs Recent models such as GPT4 (OpenAI, 2023), Codestral (Mistral AI team, 2024), and DeepSeek-Coder (Zhu et al., 2024) have advanced code generation through instruction tuning and RLHF with mixed code and natural language data (Ziegler et al., 2020; Husain et al., 2020; Rafailov et al., 2023). Code generation with LLMs has been enhanced by various techniques. Some approaches focus on improving the quality of generated code using planning algorithms, transitioning from outlines to detailed implementations (Zhang et al., 2022; Yao et al., 2023; Zelikman et al., 2023; Zhou et al., 2023; Zheng et al., 2023). Other methods sample multiple programs from the same LLM and rank them to identify the best one (Chen et al., 2023a; 2022; Ni et al., 2023). Additionally, some works leverage multi-agent collaboration frameworks to enhance code generation quality (Zhang et al., 2024; Huang et al., 2023a; Dong et al., 2024). These approaches aim to optimize the production of correct code from the outset. By contrast, MGDebugger targets the post-generation phase, focusing on debugging and fixing errors that inevitably arise during the code generation process.
Repairing LLM-Generated Code Program repair is a critical aspect of software development, aiming to automatically identify and fix bugs in code (Just et al., 2014; Gupta et al., 2020; Yasunaga & Liang, 2021). There are two main streams of research in repairing code generated by LLMs: (1) training models to repair code (Huang et al., 2023b; Jiang et al., 2024; Ding et al., 2024; Zheng et al., 2024; Moon et al., 2024; Kumar et al., 2024) and (2) providing external feedback to the raw pretrained models to fix code (Jiang et al., 2023; Chen et al., 2023b; Olausson et al., 2023; Zhong et al., 2024; Hu et al., 2024). By contrast to previous work that trains separate models for code repair (Ding et al., 2024; Zheng et al., 2024; Moon et al., 2024), MGDebugger does not require task-specific retraining but takes advantage of the inherent capabilities of pretrained LLMs. This flexibility allows MGDebugger to operate in zero-shot settings, offering a lightweight and scalable alternative. And exploring the ability of LLMs to fix their own code is a promising direction for self-improvement training of the LLMs (Wang et al., 2023; Burns et al., 2023).
MGDebugger falls under the category of work that leverages pretrained models to fix code by reasoning with external feedback. Several recent methods (Zhang et al., 2023; Olausson et al., 2023; Bouzenia et al., 2024; Lee et al., 2024; Xia & Zhang, 2023) utilize execution results from test cases to guide LLMs in code correction. More recent works have explored advanced debugging techniques utilizing LLM’s reasoning ability. Reflexion (Shinn et al., 2023) prompts LLMs to reflect on the generated code and uses a memory buffer for iterative refinement. Self-Debugging (Chen et al., 2023b) prompts LLMs to explain or dry run generated programs, known as rubber duck debugging. LDB (Zhong et al., 2024) segments programs into basic blocks, tracking variable values during runtime after each block to verify the correctness against the task description. Although these methods incorporate detailed execution feedback and iterative refinement, they treat the whole function as a single unit and perform sequential debugging, limiting their effectiveness with complex code (Xia et al., 2023; Hossain et al., 2024). MGDebugger addresses this issue by introducing a hierarchical approach, debugging from low-level errors to high-level flaws. This method ensures a more systematic and accurate debugging process, especially for complex and multifunctional systems.
3 Methodology
3.1 Overview
Figure 1: Workflow of MGDebugger compared to existing methods. Existing methods debug the function holistically, making it difficult to pinpoint the bugs. To address this issue, MGDebugger decomposes the code into a hierarchical structure, isolating subfunctions for independent bottom-up debugging. In this way, MGDebugger can identify and fix bugs at multiple levels of granularity, from bottom-level syntax errors to high-level algorithmic flaws. For simplicity, we omit the exact code after decomposition here, and provide the full example in Appendix A.
@@ 3.1 Overview
We present MGDebugger, a novel bottom-up hierarchical debugging method for repairing LLM-generated code. The overall workflow of MGDebugger is illustrated in Figure 1, while the detailed debugging process for each subfunction is depicted in Figure 2.
@@
As shown in Figure 1, MGDebugger begins with Hierarchical Code Decomposition (Section 3.2), which decomposes the input buggy code into a hierarchical structure of subfunctions. This enables systematic identification and resolution of bugs at various levels of granularity. For each subfunction, MGDebugger Generates Test Case Generation for Subfunctions (Section 3.3), deriving private test cases from public test cases of the main function, as illustrated in Figure 2. MGDebugger then executes these test cases and Debugs Subfunction with LLM-Simulated Execution (Section 3.4). The LLM simulates step-by-step code execution for failed test cases, monitoring critical variables and state changes to pinpoint the cause of errors. Once a subfunction has been fixed, MGDebugger updates it in the hierarchical structure and propagates the changes to dependent functions through Bottom-up Debugging (Section 3.5). This hierarchical debugging approach not only tackles different types of bugs at various levels of abstraction but also guarantees a cohesive and systematic debugging process throughout the entire code structure.
3.2 Hierarchical Code Decomposition
@@
Specifically, given an LLM-generated function $f$, we decompose it into a hierarchical structure of subfunctions denoted as $(f_{1},...,f_{n})$. These subfunctions can be organized as a tree $f_{\text{root}}=\text{TREE}(f_{\text{root}},\text{CHILD}(f_{\text{root}}))$, where $f_{\text{root}}$ represents the main function and $\text{CHILD}(f)$ denotes the set of subfunctions directly called by $f$. We leverage an LLM for the decomposition, adhering to three principles: (1) each subfunction represents the minimal reusable unit of code with a specific purpose, (2) higher-level functions call lower-level functions to achieve complex functionality, and (3) the overall structure facilitates isolated testing and debugging. As illustrated in Figure 1, the resulting tree-like structure allows us to isolate logical units of the code, enabling more focused debugging efforts across different levels of granularity (Woodfield et al., 1981; Isazadeh et al., 2017). The prompt template used for code decomposition is provided in Appendix G.1.
3.3 Generating Test Cases for Subfunctions
Figure 2: Illustration of the subfunction debugging process. Initially, the LLM generates test cases for the subfunction and collects the results. Subsequently, it simulates the code execution step-by-step, focusing on the change of key variables. This helps the LLM to pinpoint errors accurately and produce a corrected version of the subfunction.
@@ 3.3 Generating Test Cases for Subfunctions
Having obtained the hierarchy of subfunctions, we aim to verify the correctness of each subfunction. For this purpose, we generate test cases for each subfunction leveraging automatic unit test generation techniques (Wang et al., 2021; Schäfer et al., 2024; Liu et al., 2023a). For each subfunction $f_{i}\in f_{\text{root}}$, we generate a set of test cases $\mathcal{T}_{i}$. Following the problem settings from Chen et al. (2023b) and Zhong et al. (2024), we assume that the public test cases for the main function $\mathcal{T}_{\text{pub}}$ have been provided, which is common in most code generation benchmarks (Chen et al., 2021; Hendrycks et al., 2021; Muennighoff et al., 2023) [^2] Otherwise, we can use LLM-generated test cases instead.. We can leverage these test cases to derive a set of corresponding test cases for each subfunction.
@@
We employ the same LLM for the test case generation. For each $f_{i}\in f_{\text{root}}$. The LLM is now prompted to perform the following steps: (1) analyze how the subfunction is used within the main function and how it contributes to the expected outputs in the public test cases; (2) for each public test case, reason through the overall code structure step by step to figure out the input and expected output for the subfunction. This approach ensures that the generated test cases are not only reflective of the subfunction’s intended functionality but also contextualized within the constraints provided by the public test cases, enhancing the robustness and relevance of the test cases. The template for generating test cases is provided in Appendix G.2.
3.4 Debugging Subfunctions with LLM-simulated Execution
@@
Inspired by the methodology in Li et al. (2023), we propose an LLM-simulated code executor, which prompts LLM to act as a Python interpreter and track the code execution. As shown in Figure 2, we request the LLM to simulate the execution process, reasoning about key variables and their states at each step, and thoroughly analyzing the failed test cases. This eliminates the need for an external debugger, offering a more flexible and efficient debugging solution. In addition, the LLM can accurately identify where errors occur and grasp their surrounding context. The LLM prompt for the debugging process is detailed in Appendix G.3.
@@
6:              $f_{i}^{\prime}\leftarrow\textsc{{MGDebugger}}(f_{i},\mathcal{T}_{\text{pub}})$triangleright Recursive debugging
7:              $f_{i}$=$f_{i}^{\prime}$triangleright Replace $f_{i}$ with the debugged version
@@
10:     $\mathcal{T}_{f}\leftarrow\textsc{GenTest}(f,\mathcal{T}_{\text{pub}})$triangleright Generate test cases for $f$
11:     $\mathcal{R}_{f}\leftarrow\textsc{Exec}(f,\mathcal{T}_{f})$triangleright Execute test cases for $f$
@@
13:         return $f$triangleright Correct function; keep as is
@@
15:         $f^{\prime}\leftarrow\textsc{Debug}(f,\mathcal{T}_{f},\mathcal{R}_{f})$triangleright Debug function $f$ based on test results $\mathcal{R}_{f}$
16:         return $f^{\prime}$triangleright Return the corrected code

### v3

@@ Abstract.
While large language models have made significant strides in code generation, the pass rate of the generated code is bottlenecked on subtle errors, often requiring human intervention to pass tests, especially for complex problems. Existing LLM-based debugging systems treat generated programs as holistic units, failing to address bugs at multiple levels of granularity, from low-level syntax errors to high-level algorithmic flaws. In this paper, we introduce Multi-Granularity Debugger (MGDebugger), a hierarchical code debugger that isolates, identifies, and resolves bugs at various levels of granularity. MGDebugger decomposes problematic code into a hierarchical tree structure of subfunctions, with each level representing a particular granularity of error. During debugging, it analyzes each subfunction and iteratively resolves bugs in a bottom-up manner. To effectively test each subfunction, we propose an LLM-simulated Python executor, which traces code execution and tracks important variable states to pinpoint errors accurately. Extensive experiments with both open-source and commercial LLMs demonstrate that MGDebugger significantly outperforms existing debugging systems, achieving up to 18.9% improvement in accuracy over seed generations in HumanEval and a 97.6% repair success rate in HumanEvalFix. Furthermore, MGDebugger effectively generalizes to real-world software defects, fixing 129 bugs in Defects4J v1.2 and 64 bugs in Defects4J v2.0, outperforming state-of-the-art program repair approaches by 13.2% and 8.5% respectively. These results demonstrate MGDebugger’s robustness and effectiveness across different scenarios, establishing it as a powerful tool for closing the last mile of code generation. [^1] Code and data available at https://github.com/YerbaPage/MGDebugger
copyright: none
1. Introduction
@@
Large language models (LLMs) such as GPT-4 (OpenAI, 2023), LLaMA (Touvron et al., 2023), and DeepSeek-Coder (AI, 2024) have made significant advances in AI-assisted coding tasks (Chen et al., 2021; Li et al., 2022; Just et al., 2014). Trained on vast corpora of text and code, LLMs can understand and generate code snippets for various programming tasks, ranging from simple data structures to complex algorithmic problems (Li et al., 2022). These models have demonstrated proficiency in tasks such as code completion, bug detection, and even tackling competitive programming challenges (Chen et al., 2021; Zhang et al., 2024b; Wen et al., 2024).
@@
While the code generated by large language models generally meets the requirements, it often contains critical errors that require human intervention to pass tests (Liu et al., 2023a; Dou et al., 2024). This has gradually led to a new development paradigm: large models generate the code, while humans fix it. Therefore, the “last mile”, as well as the most crucial step, of code generation is how to efficiently repair the code generated by large models.
@@
Numerous efforts have been made to debug LLM-generated code. The most popular way is to reuse the LLM generator to debug the generated code with the feedback from test case execution (Chen et al., 2023b; Zhong et al., 2024; Hu et al., 2024). While these methods increase the pass rates, they treat the erroneous program as a holistic set of statements (i.e., as a single unit without considering the internal structure or modularity) (Chen et al., 2023b; Shinn et al., 2023; Zhong et al., 2024; Ding et al., 2024) regardless of the varying types and levels of failures. Failures of test cases arise from different levels of factors, from low-level syntactic errors to high-level algorithmic flaws. A holistic treatment overlooks the internal structure of the code and limits the effectiveness of the debugging systems, especially when dealing with complex programs that need debugging across different modules (Zeller, 2009; Tian et al., 2024).
@@
Figure 1. Workflow of MGDebugger compared to existing methods.
Extensive experiments across multiple models and benchmarks demonstrate that MGDebugger significantly outperforms existing debugging methods. Using open-source models like DeepSeek-Coder-V2-Lite, CodeQwen1.5, and Codestral, MGDebugger elevates accuracy from 75.6% to 94.5% on HumanEval (Chen et al., 2021) and achieves a remarkable 97.6% repair success rate on HumanEvalFix (Muennighoff et al., 2023). The approach also generalizes well to state-of-the-art proprietary models, improving GPT-4o and Claude 3.5 Sonnet’s performance by 6.1% on HumanEval, reaching accuracies of 96.3% and 95.7% respectively.
Our comprehensive ablation studies confirm the vital role of each component in MGDebugger, with the hierarchical debugging strategy being the most critical element. Removing this component leads to a significant drop in repair success rate from 76.3% to 52.6% on HumanEval, highlighting its importance for systematic bug identification and resolution. Detailed analyses across different bug categories reveal MGDebugger’s exceptional performance in handling various error types, achieving 100% success rates in fixing missing logic, excess logic, operator misuse, variable misuse, and function misuse bugs with DeepSeek-Coder. Furthermore, MGDebugger maintains robust performance across code snippets of varying lengths and complexities, consistently outperforming baseline methods especially for longer, more complex code where other approaches typically struggle.
While MGDebugger is primarily designed to fix LLM-generated code and close the last mile of code generation, we also evaluate its generalizability to real-world software defects. When applied to the Defects4J dataset (Just et al., 2014), MGDebugger successfully repairs 129 bugs in Defects4J v1.2 and 64 bugs in Defects4J v2.0, outperforming specialized program repair tools. This demonstrates that the hierarchical debugging principles with test case generation and LLM-simulated execution of MGDebugger extend effectively beyond LLM-generated code to human-written software with complex project structures.
The main contributions of this work can be summarized as:
  • We introduce MGDebugger, a novel hierarchical debugging method to repair LLM-generated code at multiple levels of granularity.
  • We propose an LLM-simulated execution approach that enables precise bug identification by tracing code execution and tracking variable states within subfunctions, eliminating the need for external debuggers.
  • We demonstrate MGDebugger’s effectiveness through comprehensive experiments across multiple models and benchmarks, showing significant improvements over existing methods with up to 18.9% accuracy gains on HumanEval and 97.6% repair success rate on HumanEvalFix.
  • We establish MGDebugger’s generalizability by successfully applying it to real-world software defects in Defects4J, outperforming specialized program repair tools and demonstrating its potential as a unified debugging framework for both AI-generated and human-written code.
2. Methodology
2.1. Overview
@@
We present MGDebugger, a novel bottom-up hierarchical debugging method for repairing LLM-generated code. The overall workflow of MGDebugger transforming holistic debugging into hierarchical subfunction debugging is illustrated in Figure 1, while the detailed debugging process for each subfunction is depicted in Figure 2.
@@
As shown in Figure 1, MGDebugger begins with Hierarchical Code Decomposition (Section 2.2), which decomposes the input buggy code into a hierarchical structure of subfunctions. This enables systematic identification and resolution of bugs at various levels of granularity. For each subfunction, MGDebugger Generates Test Cases for Subfunctions (Section 2.3), deriving private test cases from public test cases of the main function, as illustrated in Figure 2. MGDebugger then executes these test cases and Debugs Subfunction with LLM-Simulated Execution (Section 2.4). The LLM simulates step-by-step code execution for failed test cases, monitoring critical variables and state changes to pinpoint the cause of errors. Once a subfunction has been fixed, MGDebugger updates it in the hierarchical structure and propagates the changes to dependent functions through Bottom-up Debugging (Section 2.5). This hierarchical debugging approach not only tackles different types of bugs at various levels of abstraction but also guarantees a cohesive and systematic debugging process throughout the entire code structure.
2.2. Hierarchical Code Decomposition
@@
Specifically, given an LLM-generated function $f$, we decompose it into a hierarchical structure of subfunctions denoted as $(f_{1},...,f_{n})$. We formally define a ”subfunction” as a contiguous sequence of statements $S=[s_{1},s_{2},...,s_{k}]$ within function $f$, such that $S$ represents a logically coherent and minimal unit of functionality that can be abstracted as a function $f_{i}(x)$ with well-defined inputs and outputs. These subfunctions can be organized as a tree $f_{\text{root}}=\text{TREE}(f_{\text{root}},\text{CHILD}(f_{\text{root}}))$, where $f_{\text{root}}$ represents the main function and $\text{CHILD}(f)$ denotes the set of subfunctions directly called by $f$. We leverage an LLM for the decomposition, adhering to three principles: (1) each subfunction represents the minimal reusable unit of code with a specific purpose, (2) higher-level functions call lower-level functions to achieve complex functionality, and (3) the overall structure facilitates isolated testing and debugging. As illustrated in Figure 1, the resulting tree-like structure allows us to isolate logical units of the code, enabling more focused debugging efforts across different levels of granularity (Woodfield et al., 1981; Isazadeh et al., 2017). The prompt template for hierarchical decomposition is shown below:
Prompt Template for Hierarchical Decomposition Convert the following Python code into a tree-style hierarchical structure with multiple levels of sub-functions. Each significant step or logical block should be its own function, and functions can call other sub-functions. Ensure that the main function calls these sub-functions in the correct order, creating a tree-like structure. Original Code: {code} Instruction: Please first analyze the codes step by step, and then provide the converted code in a Python code block. When providing the final converted code, make sure to include all the functions in a flattened format, where each function is defined separately.
2.3. Generating Test Cases for Subfunctions
Figure 2. Illustration of the subfunction debugging process in MGDebugger.
@@ 2.3. Generating Test Cases for Subfunctions
Having obtained the hierarchy of subfunctions, we aim to verify the correctness of each subfunction. For this purpose, we generate test cases for each subfunction leveraging automatic unit test generation techniques (Wang et al., 2021; Schäfer et al., 2024; Liu et al., 2023b). For each subfunction $f_{i}\in f_{\text{root}}$, we generate a set of test cases $\mathcal{T}_{i}$. Following the problem settings from Chen et al. (2023b) and Zhong et al. (2024), we assume that the public test cases for the main function $\mathcal{T}_{\text{pub}}$ have been provided, which is common in most code generation benchmarks (Chen et al., 2021; Hendrycks et al., 2021; Muennighoff et al., 2023) [^2] Otherwise, we can use LLM-generated test cases instead.. We can leverage these test cases to derive a set of corresponding test cases for each subfunction.
@@
We employ the same LLM for the test case generation. For each $f_{i}\in f_{\text{root}}$. The LLM is now prompted to perform the following steps: (1) analyze how the subfunction is used within the main function and how it contributes to the expected outputs in the public test cases; (2) for each public test case, reason through the overall code structure step by step to figure out the input and expected output for the subfunction. This approach ensures that the generated test cases are not only reflective of the subfunction’s intended functionality but also contextualized within the constraints provided by the public test cases, enhancing the robustness and relevance of the test cases. The prompt template for test case generation is shown below:
Prompt Template for Test Case Generation SYSTEM_PROMPT: You are an AI assistant specialized in analyzing Python functions and generating test cases. USER_PROMPT: Full Code: {full_code} Public Test Cases for the Main Function: {public_test_cases} Instruction: Please analyze how the {function_name} function is used within the main function and how it contributes to the expected outputs in the gold test cases. For each test case, you should analyze step-by-step based on both the input and the expected output of the main function, and then provide the corresponding input and expected output for the {function_name} function. Ensure that the generated test cases are consistent with the behavior expected in the public test cases.
2.4. Debugging Subfunctions with LLM-simulated Execution
@@
Inspired by the methodology in Li et al. (2023), we propose an LLM-simulated code executor, which prompts LLM to act as a code interpreter and track the code execution. We choose LLM-simulated execution over real execution for several key reasons (La Malfa et al., 2024; Ni et al., 2024; Zheng et al., 2024): (1) Many code snippets require specific environments, hardware, or dependencies that are difficult to set up or may not be available (e.g., embedded systems, special libraries), (2) Some code cannot be executed due to compilation errors or security concerns, necessitating sandboxing or virtual execution. As shown in Figure 2, we request the LLM to simulate the execution process, reasoning about key variables and their states at each step, and thoroughly analyzing the failed test cases. This eliminates the need for an external debugger, offering a more flexible and efficient debugging solution. In addition, the LLM can accurately identify where errors occur and grasp their surrounding context. The prompt template for debugging subfunction is shown below:
Prompt Template for Debugging Subfunction SYSTEM_PROMPT: You are an AI assistant helping to debug Python functions. USER_PROMPT: Debug the following Python function. The function is not passing all test cases. Analyze the code, identify the bug, and provide a fixed version of the function. Function Code: {function_code} Test Case Results: {test_case_results} Instruction: Please try to work as a Python interpreter to execute the code step-by-step. Identify the change of each variable as you “run” the code line-by-line. Based on the execution trace, try to identify the bug and provide the final fixed code in a Python code block.
@@
6:     $f_{i}^{\prime}\leftarrow\textsc{{MGDebugger}}(f_{i},\mathcal{T}_{\text{pub}})$ $\triangleright$ Recursive debugging
7:     $f_{i}$ = $f_{i}^{\prime}$ $\triangleright$ Replace $f_{i}$ with the debugged version
@@
10:  $\mathcal{T}_{f}\leftarrow\textsc{GenTest}(f,\mathcal{T}_{\text{pub}})$ $\triangleright$ Generate test cases for $f$
11:  $\mathcal{R}_{f}\leftarrow\textsc{Exec}(f,\mathcal{T}_{f})$ $\triangleright$ Execute test cases for $f$
@@
13:   return $f$ $\triangleright$ Correct function; keep as is
@@
15:   $f^{\prime}\leftarrow\textsc{Debug}(f,\mathcal{T}_{f},\mathcal{R}_{f})$ $\triangleright$ Debug function $f$ based on test results $\mathcal{R}_{f}$
16:   return $f^{\prime}$ $\triangleright$ Return the corrected code